# Impacts of dynamic radiation transport to one-cell inversions on modern GPUs with higher-order discretization

### v1

@@ 2.5 Implementation on GPUs
Implementing the OCI and SI approaches on GPUs requires a numerical linear algebra solver library like LAPACK [20]. Many high-performance open-source linear algebra tools exist (e.g., Trillinos, PETSc, SUNDIALS, MAGMA), but we chose a vendor-supplied package depending on the hardware target of choice. Our target hardware is an AMD MI250X so we use the AMD ROCm compute library to solve the system of equations. Modern GPU vendor-supplied LAPACK libraries often include a batched class of solvers, which operate on a group of like-sized systems in unison and are optimized by the hardware vendors. For example, LU decomposition with pivoting (a generic direct solver for a system of linear equations) used in this work comes from RocSolver’s strided_batched_dgesv [21].
We use direct solvers here because all systems are “small”, with orders ranging between 4 and 100. This makes the use of a batched implementation of LU decomposition with pivoting ideal. Furthermore LAPACK-type implementations of _gesv automatically return the $L+U+D$ decomposition of $A_{j}$. So, in subsequent iterations, this system can be back solved quickly (using LAPACK _getrs). In this mode for both SI and OCI, the only user-defined device kernels are the RHS vector builders which are already memory safe operations.
This software engineering design will increase the memory footprint of OCI and SI as the $A_{j}$ matrices are stored in memory. This is acceptable for 1D transport but more optimization may be required when moving to 2D and 3D solvers.
@@
build $A$ in all cells and move to Device
@@
      build constant part of $b$ in all cells
      move constant part of $b$ to Device
@@
        //transport sweep
@@
            build variable part of $b$ at cells $j$ and $J-i$
@@
                    //in Aj out L+U+D
                   $\Psi_{j}$ = GPU_strided_batched_dgesv($A$[$\beta^{2}j$],$b$[$\beta j$])
@@
                   $\Psi_{j}$ = GPU_strided_batched_dgetrs($A$[$\beta^{2}j$],$b$[$\beta j$])
@@
      communicate group to group sources
@@
Algorithm 1 Source iteration algorithm implemented on GPU. Simplified for brevity.
@@
build $A$ in all cells and move to device
build constant part of $b$ in all cells and move to device
@@
OCI’s systems are represented as dense within a cell and built in a strided-batched configuration to take advantage of the block sparsity. However, now systems within an iteration can be dispatched in unison. The intra-iteration $b$-vector production kernels are the only user-defined device functions required in this algorithm. These are relatively simple to implement as they are thread-safe operations.
@@ 4 Discussion
Parallel sweeping algorithms may not be well suited to GPUs where non-uniform work distributions come with significant overhead. Available results for full parallel sweeps on GPUs show that even optimized applications underperform relative to the theoretical hardware resources available [23, 24, 25]. On the other hand, space-parallel OCI uses the same parallel scheme in 2D and 3D as it does in 1D, with arithmetically intense operations that align well with the GPU parallelism paradigm. So, we hypothesize that OCI can better take advantage of the compute resources available on GPUs in higher dimensions than full-parallel-sweep SI on GPUs, but this requires further study.
@@ 5 Conclusions and Future Work
Moving forward, we are exploring synthetic acceleration techniques to preserve the OCI space-parallel performance on GPUs while ameliorating issues in the thin limit. Space-parallel OCI schemes offer promise as a high-performing class of iterative solvers for time-dependent radiation transport on modern heterogeneous compute architectures.

### v2

@@ 2.5 Implementation on GPUs
Implementing the OCI and SI approaches on GPUs requires a numerical linear algebra solver library like LAPACK [20]. Many high-performance open-source linear algebra tools exist (e.g., Trillinos [21], PETSc [22], MAGMA [23]), but we chose a vendor-supplied package depending on the hardware target of choice. Our target hardware is an AMD MI250X so we use the AMD ROCm compute library to solve the system of equations. Modern GPU vendor-supplied LAPACK libraries often include a batched class of solvers, which operate on a group of like-sized systems in unison and are optimized by the hardware vendors. For example, LU decomposition with pivoting (a generic direct solver for a system of linear equations) used in this work, comes from RocSolver’s strided_batched_dgesv [24].
We use direct solvers here because all systems are relatively small, with orders ranging between 4 and 100. This makes the use of a batched implementation of LU decomposition with pivoting ideal. Furthermore, LAPACK-type implementations of _gesv (GEneral linear SolVe) automatically return the $L+U+D$ decomposition of a generic matrix $A$. So, in subsequent iterations, this system can be back solved quickly (using LAPACK _getrs). In this mode for both SI and OCI, the only user-defined device kernels are the RHS vector builders which are already memory safe operations.
This software engineering design will increase the memory footprint of OCI and SI as the $LHS$ matrices are stored in memory. This is acceptable for 1D transport but more optimization may be required when moving to 2D and 3D solvers.
@@
build $\mathbf{L}_{c,j,g,m}$ for each cell, angle, and group
  //Eq. 42
move $\mathbf{L}_{c,j,g,m}$ to device
@@
      build constant part of $\bm{d}_{j,g,m}$ in all cells, angles, and groups
        //Eq. (43)
      move constant part of $\bm{d}_{j,g,m}$ to device
@@
        //parallel over groups and angles
@@
            build variable part of $\bm{d}_{j,g,m}$ at cells $j$ and $J-j$
              //on GPU Eq. (43)
@@
                    //LHS in, L+U+D out
                   $\Psi_{j}$ = GPU_strided_batched_dgesv($\mathbf{L}_{c,j}$[$\beta^{2}j$],$\bm{d}_{j}$[$\beta j$])
@@
                   $\Psi_{j}$ = GPU_strided_batched_dgetrs($\mathbf{L}_{c,j}$[$\beta^{2}j$],$\bm{d}_{j}$[$\beta j$])
@@
      communicate group to group
        //Eq. (45)
@@
Algorithm 1 Source iteration algorithm implemented on GPU where $\bm{\phi}$ is scalar flux. Equations in Appendix A. Simplified for brevity.
@@
build $\bm{A}_{j}$ in all cells and move to device
  //Eq. (40b)
build constant part of $\bm{b}_{j}$ in all cells and move to device
  //Eq. (40c)
@@
        //incident angular fluxes from previous iteration
        //user-defined GPU kernel
       build variable part of $\bm{b}_{j}$ in all cells
        //Eq. (40c)
@@
OCI’s systems are represented as dense within a cell and built in a strided-batched configuration to take advantage of the block sparsity. However, now systems within an iteration can be dispatched in unison. Just as with the SI algorithm, the GPU strided batched solver implements the parallel loop over all cells. The intra-iteration $b$-vector production kernels are the only user-defined device functions required in this algorithm. These are relatively simple to implement as they are thread-safe operations.
@@ 4 Discussion
Parallel sweeping algorithms may not be well suited to GPUs where non-uniform work distributions come with significant overhead. Available results for full parallel sweeps on GPUs show that even optimized applications underperform relative to the theoretical hardware resources available [26, 27, 28, 29, 30]. On the other hand, space-parallel OCI uses the same parallel scheme in 2D and 3D as it does in 1D, with arithmetically intense operations that align well with the GPU parallelism paradigm. So, we hypothesize that OCI can better take advantage of the compute resources available on GPUs in higher dimensions than full-parallel-sweep SI on GPUs, but this requires further study.
Production SI codes that simulate high-fidelity 3D models never form the actual iteration matrix, resulting in significant memory efficiency [31, 32, 33, 34, 35, 36]. Our implementation of OCI is not memory efficient, because we store the strided batched representation of the entire linear system for the iteration algorithm. While this block-sparse representation does dramatically save memory, more optimizations to OCI may be needed when moving to 3D. For example, OCI forms the iteration matrix as a linear operator. Many algorithms exist to solve systems of linear operators without ever forming the matrix. Libraries already exist to do this on CPUs and GPUs (e.g., LibCEED [37]). Even more simple optimization may involve using sparse representations for the cell blocks themselves and sparse direct solvers.
OCI algorithms may be better suited for domain-decomposed simulations (required for distributed-memory parallelism) than SI. Some KBA algorithms can degrade with large processor counts [5]. For a hypothetical OCI algorithm, communication between subdomains can be very organized, occurring at every iteration (or after a specified number of iterations) at the same time as all other subdomains. This may lead to superior weak scaling across large decomposed problems on distributed (MPI) type systems. Many parallel domain decomposition algorithms are already based on the same underlying parallel block Jacobi, Gauss–Seidel, or red-black iterations as OCI [38, 25].
While this work is limited to AMD GPUs and the ROCm software stack, our implementation of OCI is portable to other GPU types. In fact, the workflow present in OCI—solving many small dense or sparse linear algebra systems in parallel—is common in much numerical software. We expect that these types of solvers will exist when porting to a new hardware accelerator used for scientific computing. For example, Nvidia’s cuSOLVER [39] and Intel’s oneMKL [40] both implement strided batched versions of LU decomposition with pivoting for their GPUs. Porting the OCI algorithm presented in Algorithm 2 to another GPU would just require altering names of solver function calls, compiler commands, and switching out library header imports.
@@ 5 Conclusions and Future Work
Moving forward, we are exploring synthetic acceleration techniques to preserve the OCI space-parallel performance on GPUs while ameliorating issues in the thin and scattering limits. Space-parallel OCI schemes offer promise as a high-performing class of iterative solvers for time-dependent radiation transport on modern heterogeneous compute architectures.
@@ Acknowledgments
Appendix A Source Iteration Systems
Using a source iteration to solve a multi-group problem with multiple balance in time [15] and simple corner balance in space [2] gives a $4\times 4$ system of linear equations:
$\mathbf{L}_{c,j,g,m}\begin{bmatrix}\psi_{m,k,g,j,L}^{(l+1/2)}\\ \psi_{m,k,g,j,R}^{(l+1/2)}\\ \psi_{m,k+1/2,g,j,L}^{(l+1/2)}\\ \psi_{m,k+1/2,g,j,R}^{(l+1/2)}\end{bmatrix}=\mathbf{d}_{j,m,g}\;,$ (42)
solved in every angle and group by sweeping from cell to cell, where $\mathbf{L}_{c,j,g,m}$ is from Eq. (8b) and
$\mathbf{d}_{j,g,m}=\begin{cases}\bm{d}_{j,g,m}^{+}&\mu_{m}>0\\ \bm{d}_{j,g,m}^{-}&\mu_{m}<0\\ \end{cases}\;,$ (43a)
where
$\bm{d}_{j,g,m}^{+}=\begin{bmatrix}\frac{\Delta x_{j}}{4}\left(\Sigma_{s,g\rightarrow g,j}\phi_{k,g,j,L}^{(l)}+Q_{k,j,L}\right)+\frac{\Delta x_{j}}{2v_{g}\Delta t}\psi_{m,k-1/2,g,j,L}+\mu_{m}\psi_{m,k,g,j-1,R}^{(l+1/2)}\\ \frac{\Delta x_{j}}{4}\left(\Sigma_{s,g\rightarrow g,j}\phi_{k,g,j,R}^{(l)}+Q_{k,j,R}\right)+\frac{\Delta x_{j}}{2v_{g}\Delta t}\psi_{m,k-1/2,g,j,R}\\ \frac{\Delta x_{j}}{4}\left(\Sigma_{s,g\rightarrow g,j}\phi_{k+1/2,g,j,L}^{(l)}+Q_{k+1/2,j,L}\right)+\mu_{m}\psi_{m,k+1/2,g,j-1,R}^{(l+1/2)}\\ \frac{\Delta x_{j}}{4}\left(\Sigma_{s,g\rightarrow g,j}\phi_{k+1/2,g,j,R}^{(l)}+Q_{k+1/2,j,R}\right)\end{bmatrix}\;,$ (43b)
$\bm{d}_{j,g,m}^{-}=\begin{bmatrix}\frac{\Delta x_{j}}{4}\left(\Sigma_{s,g\rightarrow g,j}\phi_{k,g,j,L}^{(l)}+Q_{k,j,L}\right)+\frac{\Delta x_{j}}{2v_{g}\Delta t}\psi_{m,k-1/2,g,j,L}\\ \frac{\Delta x_{j}}{4}\left(\Sigma_{s,g\rightarrow g,j}\phi_{k,g,j,R}^{(l)}+Q_{k,j,R}\right)+\frac{\Delta x_{j}}{2v_{g}\Delta t}\psi_{m,k-1/2,g,j,R}-\mu_{m}\psi_{m,k,g,j+1,L}^{(l+1/2)}\\ \frac{\Delta x_{j}}{4}\left(\Sigma_{s,g\rightarrow g,j}\phi_{k+1/2,g,j,L}^{(l)}+Q_{k+1/2,j,L}\right)\\ \frac{\Delta x_{j}}{4}\left(\Sigma_{s,g\rightarrow g,j}\phi_{k+1/2,g,j,R}^{(l)}+Q_{k+1/2,j,R}\right)-\mu_{m}\psi_{m,k+1/2,g,j+1,L}^{(l+1/2)}\end{bmatrix}\;,$ (43c)
and $\phi$ is the scalar flux.
After sweeping the mesh cells in the appropriate directions for each angle in the quadrature set and every group, the scalar flux vector can be updated via
$\bm{\phi}_{k,g,j}^{(l+1/2)}=\begin{bmatrix}\phi_{k,g,j,L}^{(l+1/2)}\\ \phi_{k,g,j,R}^{(l+1/2)}\\ \phi_{k+1/2,g,j,L}^{(l+1/2)}\\ \phi_{k+1/2,g,j,R}^{(l+1/2)}\end{bmatrix}=\sum\limits_{n=1}^{N}w_{n}\begin{bmatrix}\psi_{n,k,g,j,L}^{(l+1/2)}\\ \psi_{n,k,g,j,R}^{(l+1/2)}\\ \psi_{n,k+1/2,g,j,L}^{(l+1/2)}\\ \psi_{n,k+1/2,g,j,R}^{(l+1/2)}\end{bmatrix}\;.$ (44)
Then, group-to-group communication can be computed with
$\bm{\phi}_{g}^{(l+1)}=\bm{\phi}_{g}^{(l+1/2)}+\sum_{g^{\prime}\neq g}\Sigma_{s,g^{\prime}\rightarrow g}\bm{\phi}_{g^{\prime}}^{(l+1/2)}\;.$ (45)
Note that within group scattering is computed in the transport sweep itself, in Eq. (43). This algorithm allows for all groups and angles to be solved in parallel (using a Jacobi iteration). This algorithm gives source iterations the greatest number of degrees of freedom to parallelize for a 1D, slab geometry, multi-group problem—the best case scenario for GPU performance.
Then, the source iteration can continue until
$||\bm{\phi}^{(l+1)}-\bm{\phi}^{(l)}||_{2}<\epsilon(1-\rho_{e,SI})\;,$ (46a)
where $\epsilon$ is the convergence tolerance and
$\rho_{e,SI}=\frac{||\bm{\phi}^{(l+1)}-\bm{\phi}^{(l)}||_{2}}{||\bm{\phi}^{(l)}-\bm{\phi}^{(l-1)}||_{2}}\;,$ (46b)
is an empirical estimation of the spectral radius computed at every iteration of a transport solve. After converging, the simulation can move to the next time step.